A local bank report that 80% of its customers maintain a checking account 60% have a savings account and 50% have both. If a customer is chosen at random, what is the probability the customer does not have either a checking or a saving account.? What is the probability the customers does not have either a checkings or savings account?
Answer by stanbon(75887)

A local bank report that 80% of its customers maintain a checking account 60% have a savings account and 50% have both. If a customer is chosen at random, what is the probability the customer does not have either a checking or a saving account.? What is the probability the customers does not have either a checkings or savings account?
-----------
P(check OR savings) = 0.6+0.5-0.8 = 0.3
---
Ans: P(c OR s)' = 1 - 0.3 = 0.7
------------------------------------
Cheers,
Stan H.
------------------